# From Hypothesis to Publication: A Comprehensive Survey of AI-Driven Research Support Systems

### v3

@@ 7.1 Integration of Diverse Research Tasks
The research process is an integrated pipeline of interdependent stages. Paper recommendation and literature review provide an AI tool with a field overview and relevant works, ensuring that hypothesis generation is informed and of higher quality. Hypothesis validation assesses feasibility both logically and practically, with results feeding back to refine the hypothesis. In manuscript writing, validated hypotheses and prior outputs serve as key inputs. Peer review evaluates the manuscript and offers feedback across modules, enabling the hypothesis generator to adjust content accordingly (Lu et al., 2024). In addition, combinations can also be made between some small fields, for instance, meta-review generation could be integrated with scientific claim verification, experiment verification could be linked with hypothesis formulation (Jansen et al., 2025; Yuan et al., 2025; Liu et al., 2024d), and research paper recommendation systems could be connected with manuscript writing processes (Gu and Hahnloser, 2023). Furthermore, some studies have begun to emphasize the development of systems capable of covering multiple stages of the research process (Jansen et al., 2024; Weng et al., 2024; Yu et al., 2024a).
@@ 7.2 Integration with Reasoning-Oriented Language Models
Research is a process that places a significant emphasis on logic and reasoning. Theorem proving serves as a subtask within logical reasoning (Li et al., 2024e), while hypothesis generation is widely recognized as the primary form of reasoning employed by scientists when observing the world and proposing hypothesis to explain these observations (Yang et al., 2024b). Experiment verification, in turn, demands a high degree of planning capability from models (Kambhampati et al., 2024). Recent advances in reasoning-oriented language models, such as OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025), have substantially enhanced the reasoning abilities of these models. Consequently, we posit that integrating reasoning language models with reasoning tasks is a promising future direction. This prediction was validated by experiments conducted by  Schmidgall et al. (2025) using o1-Preview.
@@ 8 Conclusion
We have made our best effort, but there may still be some limitations. Due to space constraints, we provide only concise summaries of each method without detailed technical elaboration. Given the rapid progress in AI and the expanding research landscape, we primarily focus on works published after 2022, with earlier studies receiving less attention. To emphasize areas that closely mimic the human research process, some topics are excluded from the main text but briefly discussed in Appendix §A. Moreover, as AI for Research is still an emerging field, the lack of standardized benchmarks and evaluation metrics hinders direct comparison. Nonetheless, we offer a comparative analysis of representative methods across domains using attribute graphs in Appendix §D.
Appendix A Further Discussion
We posit that AI for research constitutes a subset of AI for science. While AI for research primarily focuses on supporting or automating the research process, it is not domain-specific and places greater emphasis on methodological advancements. In contrast, AI for science extends beyond the research process to include result-oriented discovery processes within specific domains, such as materials design, drug discovery, biology, and the solution of partial differential equations (Zheng et al., 2023b; AI4Science and Quantum, 2023; Zhang et al., 2024d).
Hypothesis generation, which is primarily based on literature-based review (LBD) (Swanson, 1986; Sebastian et al., 2017), emphasizing the process by which researchers generate new concepts, solutions, or approaches through existing research and their own reasoning. Scientific discovery encompasses not only hypothesis generation, but also innovation in fields like molecular optimization and drug development  (Ye et al., 2024; Liu et al., 2024b), driven by outcome-oriented results.
Existing research frequently addresses the systematic literature survey, which constitutes a component of the knowledge synthesis process during hypothesis formulation, alongside the related work generation phase in manuscript writing (Luo et al., 2025). However, we argue that these two tasks are distinct in nature. The systematic literature survey primarily focuses on summarizing knowledge extracted from diverse scientific documents, thereby assisting researchers in acquiring an initial understanding of a specific field (Altmami and Menai, 2022). In contrast, related work generation focuses on the writing process, emphasizing selection of pertinent literature and effective content structuring  (Nishimura et al., 2024).
  • In research paper recommendation, PaperWeaver (Lee et al., 2024) offers an interactive page that allows users to modify the topics they are interested in.
  • In systematic literature review,  Block and Kuckertz (2024) highlights the significant role of humans, including setting correct questions and individualized problem-solving and theorizing. Meanwhile,  Hsu et al. (2024) emphasizes manual correction during the outline generation process.
  • In hypothesis generation, AI engages more closely with human researchers, ranging from scenarios where humans provide the core ideas and AI contributes by iteratively refining them  (Pu et al., 2024), to more collaborative settings where humans and AI engage in dialogue to facilitate new scientific discoveries  (Ni et al., 2024; Liu et al., 2024b; Ye et al., 2024).
  • In scientific claim verification,  Altuncu et al. (2023); Das et al. (2023) highlight the critical role of experts in countering fake scientific news and advocate for the incorporation of expert opinions as a form of evidence.
  • In theorem proving,  Song et al. (2024) proposes leveraging LLMs as assistants to human researchers by generating suggested proof steps throughout the proving process.
  • In experiment Validation,  Ni et al. (2024) enhances the experimental setup through human-AI dialogue, whereas  Li et al. (2024d) incorporates human input and real-time adjustments during the execution phase to optimize experimental design.
  • In manuscript writing,  Ifargan et al. (2024); Feng et al. (2024); Du et al. (2022a) require human intervention to suggest improvements to AI-generated paragraphs and enhance their quality through interactive methods.
  • In peer review,  Kumar et al. (2023); Darrin et al. (2024) advocate for assigning the responsibility of generating meta-reviews to human researchers. The role of AI is to assist by identifying conflicts among reviewers’ opinions and supporting the chair in the scoring process, rather than independently assigning scores.
Discussion: The involvement of AI in manuscript writing
The application of AI in manuscript writing has been accompanied by significant controversy. As LLMs demonstrated advanced capabilities, an increasing number of researchers began adopting these systems for scholarly composition  (Liang et al., 2024b; Gao et al., 2023a). This trend raised concerns within the academic community  (Salvagno et al., 2023), with scholars explicitly opposing the attribution of authorship to AI systems  (Lee, 2023). Despite these reservations, the substantial time efficiencies offered by this technology led researchers to gradually accept AI-assisted writing practices  (Gruda, 2024; Huang and Tan, 2023; Chen, 2023).This shift ultimately led to formal guidelines issued by leading academic journals  (Ganjavi et al., 2024; Xu, 2025).
Discussion: Some areas that have not been discussed
In addition to the eight areas discussed above, there are other lines of work that also aim to support scientific research, such as reading assistance (Kang et al., 2020; Head et al., 2021; Lo et al., 2023), which helps researchers read academic papers; literature processing [^2] https://sdproc.org, which handles documents in various formats to provide effective data for subsequent tasks; as well as code and data generation (Bauer et al., 2024; Zheng et al., 2023a), which serve as a foundation for experimental validation. However, as our focus is on the core process of scientific research, we have chosen to omit these aspects from the main text.
Appendix B Challenges
B.1 Hypothesis Formulation
Knowledge Synthesize
Existing paper recommendation tools predominantly rely on the metadata of existing publications to suggest related articles, which often results in a lack of user-specific targeting and insufficiently detailed presentation that hampers comprehension. Leveraging LLMs can facilitate the construction of dynamic user profiles, enabling personalized literature recommendations and enhancing the richness of the contextual information provided for each recommended article, ultimately improving the user experience. In the process of generating systematic literature reviews, our practical experience reveals that the outline generation tools often produces redundant results with insufficient hierarchical structure. Moreover, the full-text generation process is prone to hallucinations—for instance, statements may not correspond to the cited articles—a pervasive issue in large language models (Huang et al., 2023; Bolaños et al., 2024; Susnjak et al., 2024). This problem can be ameliorated by enhancing the foundational model capabilities or by incorporating citation tracing.
Hypothesis Generation
Most existing tools generate hypotheses by designing prompts or constructing systematic frameworks, which heavily rely on the capabilities of pre-trained models. However, these methods struggle to balance the novelty, feasibility, and validity of the hypotheses (Li et al., 2024c).Furthermore, our investigation reveals that many current approaches adopt novelty and feasibility as evaluation metrics; these metrics are either difficult to quantify or require manual scoring, which can introduce bias. To date, there is no unified benchmark to compare the various methods, and we believe that future research should prioritize establishing a unified metric that objectively reflects the strengths and weaknesses of different approaches.
B.2 Hypothesis Validation
Most existing scientific claim verification tools are largely confined to specific domains, exhibiting poor generalizability, which limits their practical applicability  (Vladika and Matthes, 2023). Theorem proving, the scarcity of relevant data adversely affects performance improvements through training , results across different proof assistants are not directly comparable, and the lack of standardized evaluation benchmarks presents numerous challenges. Moreover, current approaches remain predominantly in the research stage and lack practical tools that facilitate interaction with researchers  (Li et al., 2024e). Experiment Validation, as automatically generated experiments often suffer from a lack of methodological rigor, practical feasibility, and alignment with the original research objectives  (Lou et al., 2024). All these fields require rigorous logical reasoning, and I believe that the recent surge in advanced reasoning technologies could potentially address these issues.
B.3 Manuscript Publication
Similar to systematic literature surveys, manuscript writing is also adversely affected by hallucination issues (Athaluri et al., 2023; Huang et al., 2023). Even when forced citation generation is employed, incorrect references may still be introduced (Aljamaan et al., 2024). Furthermore, the text generated by models requires meticulous examination by researchers to avoid ethical concerns, such as plagiarism risks (Salvagno et al., 2023). AI-generated manuscript reviews frequently provide vague suggestions and are susceptible to biases (Chamoun et al., 2024; Drori and Te’eni, 2024). Additionally, during meta-review generation, models can be misled by erroneous information arising from the manuscript review process (Kumar et al., 2023). To address these issues, it may be necessary for the industry to establish appropriate regulations or to employ AI-based methods for detecting AI-generated papers and reviews (Lin et al., 2023a).
Appendix C Ethical Considerations
AI has demonstrated significant potential in enhancing productivity by mitigating human limitations, thereby motivating increased investigation into its capacity to accelerate the research process  (Messeri and Crockett, 2024). Nevertheless, its integration into scientific workflows introduces a range of ethical concerns  (Fecher et al., 2023; Morris, 2023), including algorithmic biases, data privacy issues, risks of plagiarism, and the broader implications of AI-generated content for research communities. In this work, we examine these ethical challenges across the key stages of the research lifecycle: hypothesis formulation, validation, and publication.
During the hypothesis formulation stage, research paper recommendation systems and literature reviews are commonly employed; however, they often suffer from limitations that can lead to the formation of information bubbles and restrict exposure to diverse viewpoints. Furthermore, these systems tend to reinforce recognition disparities between prominent and lesser-known researchers and may inadvertently contribute to the dissemination of misinformation  (Polonioli, 2021; Bolaños et al., 2024). To address these biases, recommendation algorithms can be enhanced by emphasizing content-based rather than author-based recommendations and by incorporating robust evaluation mechanisms to strengthen the credibility of suggested materials.
In contrast, AI-driven hypothesis generation presents more pronounced ethical challenges. First, the attribution of intellectual property rights and authorship for AI-generated hypotheses remains ambiguous  (Majumder et al., 2024a). Additionally, the widespread generation of low-quality content poses a risk of diluting the integrity of the academic landscape (Hu et al., 2024a), while the potential misuse of such technologies for illicit purposes cannot be overlooked  (Si et al., 2024). Addressing these concerns necessitates the development of robust accountability frameworks, the assignment of clear responsibility for AI-generated outputs to researchers, and the establishment of appropriate legal and regulatory mechanisms.
During the hypothesis validation phase, automated systems for scientific fact-checking remain underdeveloped. This limitation may be exploited by malicious actors to create advanced misinformation generators capable of circumventing existing fact-checking tools  (Wadden et al., 2022b). Likewise, in the context of experimental validation, there is a risk of unethical or legally questionable experiments being designed  (Eger et al., 2025). These concerns underscore the need for continued research into model safety.
During the manuscript publication stage, several challenges remain. Text generated by AI models may carry a risk of plagiarism  (Salvagno et al., 2023; Gupta and Pruthi, 2025), while AI-assisted peer reviews often offer vague feedback and exhibit inherent biases  (Schintler et al., 2023; Drori and Te’eni, 2024; Pataranutaporn et al., 2025). To address these issues, the development of robust detection methods is essential. However, current detection tools are still in the early stages of maturity  (Gupta and Pruthi, 2025).
Appendix D Ability Comparison
An effective survey should not only summarize existing methods within a field but also provide comparative analyses of different approaches. However, the domain of AI for Research remains in its early stages, with many areas lacking standardized benchmarks and even established evaluation metrics. To facilitate a clearer understanding of the distinctions among various methods, we draw on existing literature (Kang et al., 2023; Bolaños et al., 2024; Luo et al., 2025; Vladika and Matthes, 2023; Yang et al., 2023c; Li and Ouyang, 2022, 2024; Lin et al., 2023a) and adopt attribute graphs to compare representative approaches within each subfield, as illustrated in table  §1 to table  §8.
ComLittee (Kang et al., 2023) - Authorship Graphs Meta data with relevant authors R, Co, Ci
ArZiGo (Pinedo et al., 2024) - User Interest Meta data R
PaperWeaver (Lee et al., 2024) Collected Papers Meta data with description R
Kang et al. (2022) - - Author’s social network relationships +Reference relationship Meta data with relevant authors R
Table 1: Research Paper Recommendation, we referred to  Kang et al. (2023) for comparing different methods, where R represents Paper recommender score, Co represents Co-author relationship, and Ci represents Cited author relationship.
AutoSurvey (Wang et al., 2024e) Any - Outline Generation, +Full-text Generation Title & Full Content Literature Survey LLM & Human
CHIME (Hsu et al., 2024) Biomedicine - Outline Generation Title & Full Content Hierarchical Outline Automatic Metrics
Knowledge Navigator (Katz et al., 2024) Any - - Outline Generation Title & Full Content Hierarchical Outline LLM & Human
Relatedly (Palani et al., 2023) Any - - Full-text Generation Title & Related Work Literature Survey Human
STORM (Shao et al., 2024) Any - - Outline Generation, +Full-text Generation Title & Full Content Literature Survey LLM & Automatic Metrics
Table 2: Scientific Literature Review, we referred to Bolaños et al. (2024) and made modifications, thereby comparing different methods.
COI (Li et al., 2024a) Any - - -
Learn2Gen (Li et al., 2024c) Artification Intelligence - - -
MatPilot (Ni et al., 2024) Materials Science - - -
SciAgents (Ghafarollahi and Buehler, 2024) Any - - - -
SciMON (Wang et al., 2024c) Any - - - - - -
Table 3: Hypothesis Generation, we referred to Luo et al. (2025) and made modifications, thereby comparing different methods.
MULTIVERS (Wadden et al., 2022b) Claim & scientific abstract Provided - Longformer Document Label & sentence-level rationales
SFAVEL (Bazaga et al., 2024) Claim Pre-trained Language Model - - knowledge graph Top-K Facts & Corresponding Relevance Scores
ProToCo (Zeng and Gao, 2023) Claim-Evidence Pair Provided - - Sentence Label
MAGIC (Kao and Yen, 2024) Claim Provided - Dense Passage Retriever Sentence Label
aedFaCT (Altuncu et al., 2023) News Article Google Search Human Document Evidence
Table 4: Scientific Claim Verification, we referred to Vladika and Matthes (2023) and made modifications, thereby comparing different methods.
IBR (Qu et al., 2022) - Informal -
GPT-f (Polu and Sutskever, 2020) - Formal
DT-Solver (Wang et al., 2023b) Formal
POETRY (Wang et al., 2024a) - - Formal
Table 5: Theorem proving, we referred to Yang et al. (2023c) and made modifications, thereby comparing different methods.
AutoML-GPT (Zhang et al., 2023) Artification Intelligence - - - Automated Machine Learning Task-oriented Prompts -
Chemcrow (Bran et al., 2024) Chemistry - - Chemical Task Task Description
DOLPHIN (Yuan et al., 2025) Any - Automated Scientific Research Idea
MechAgents (Ni and Buehler, 2023) Physics - - Mechanical Problem - -
Manning et al. (2024) Social Science - Simulating Human - -
Method Across Stages Human Interaction Task Input Evaluation Method
AI Scientist (Lu et al., 2024) - Full-text Generation Manuscript Template & Experimental Results & Hypothesis LLM
data-to-paper (Ifargan et al., 2024) Full-text Generation Experimental Results & Hypothesis -
ScholaCite (Martin-Boyle et al., 2024) - - Related Work Generation Title & Abstract & Citation Citation Graph Metrics
SciLit (Gu and Hahnloser, 2023) - Citation Generation Keywords Automatic Metrics
Gu and Hahnloser (2024) - - Citation Generation Citation Intent & Keywords Human
Table 7: Manuscript Writing, we referred to Li and Ouyang (2022, 2024) and made modifications, thereby comparing different methods.
Gamma-Trans (Muangkammuen et al., 2022) - - - - Peer-review Score
MARG (D’Arcy et al., 2024a) - - - - Peer-review Comments
CycleResearcher (Weng et al., 2024) - - - Peer-review Comments & Score
PeerArg (Sukpanichnant et al., 2024) - - - - - Final Decision
GLIMPSE (Darrin et al., 2024) - - - - Summary of Peer-review
Table 8: Peer Review, we referred to Lin et al. (2023a) and made modifications, thereby comparing different methods.
Task Benchmark Domain Size Input Output Metric
SCHOLAT Li et al. (2020) Research Paper Recommendation 34,518 - - Recall, Precission, F1-score
ACL selection network Tao et al. (2020) Research Paper Recommendation 18,718 Topics Related Papers Accuracy
CiteSeer Kang et al. (2021) Research Paper Recommendation 1,100 Paper Related Papers Correlation Coefficient
SciReviewGen Kasanishi et al. (2023) Systematic Literature Review 10,000+ Abstracts literature review ROUGE
FacetSum Meng et al. (2021) Systematic Literature Review 60,024 Source Text+Facet Summary of Facet ROUGE
BigSurvey Liu et al. (2022) Systematic Literature Review 7,000+ Abstracts Survey Paragraph ROUGE, F1-score
SCHOLARQABENCH Asai et al. (2024) Systematic Literature Review 2,200 Question Answer with Citations Accuracy, Coverage, Citations + Relevance, Usefulness
HiCaD Zhu et al. (2023) Systematic Literature Review 7,600 Reference Papers Catalogues Catalogue Edit Distance Similarity (CEDS) + Catalogue Quality Estimate (CQE)
CLUSTREC-COVID Katz et al. (2024) Systematic Literature Review 2,284 Titles, Abstracts Topic Clusters per Topic
CHIME Hsu et al. (2024) Systematic Literature Review 2,174 Topic Hierarchies F1-score
Tian et al. (2024) Systematic Literature Review 700 Subject, Reference Title,Content -
MASSW Zhang et al. (2024c) Hypothesis Generation 152000 Context of Literature Hypothesis BLEU, ROUGE, BERTScore, + Cosine Similarity, BLEURT
IdeaBench Guo et al. (2024) Hypothesis Generation 2,374 Instruction, Background Information Hypothesis Insight Score, BERTScore, Novelty, + LLM Similarity Rating, Feasibility
SCIMON Wang et al. (2024c) Hypothesis Generation - Background Context Idea ROUGE, BERTScore +BARTScore, Novelty
MOOSEYang et al. (2024b) Hypothesis Generation 50 Background, Inspiration Hypothesis Validness, Novelty + Helpfulness
DISCOVERYBENCH Majumder et al. (2024b) Hypothesis Generation 1,167 Data Discovery Hypothesis Match Score
Hypothesis LiveIdeaBench Ruan et al. (2024a) Hypothesis Generation - Scientific Keywords Idea Originality, Feasibility + Fluency, Flexibilit
Formulation Kumar et al. (2024) Hypothesis Generation 100 Paper without Future Work Idea Idea Alignment Score, Idea Distinctness Index
SciRIFF Wadden et al. (2024) Scientific Claim Verification 137,000 Evidence, Task prompt Structured Paragraph F1, BLEU
SCIFACT Wadden et al. (2020) Scientific Claim Verification 1,409 Claim, Evidence Rationale Sentences, Label Precision, Recall, Micro-F1
SCIFACT-OPEN Wadden et al. (2022a) Scientific Claim Verification 279 Claim, Evidence Rationale Sentences, Label Precision, Recall,Micro-F1
MISSCI Glockner et al. (2024b) Scientific Claim Verification 435 Claim, Premise, Context Verification Micro F1-score,P@1,Arg@1 + METEOR Score,BERTScore +NLI-A, NLI-S, Matches@1
FEVER Thorne et al. (2018) Scientific Claim Verification 185,445 Claim, Evidence Label, Necessary Evidence F1-Score,Oracle Accuracy + Accuracy,Recall
XClaimCheck Kao and Yen (2024) Scientific Claim Verification 16,177 Claim, Evidence Label, Argument Macro-F1, Accuracy
HEALTHVER Sarrouti et al. (2021) Scientific Claim Verification 14330 Claim, Evidence Label Macro Precision, Macro Recall + Macro F1-score, Accuracy
QuanTemp V et al. (2024) Scientific Claim Verification 15,514 Claim, Evidence Label Weighted-F1 Score, Macro-F1, BLEU, + BERTScore, Cohen’s Kappa Score + Human Evaluation
SCITAB Lu et al. (2023) Scientific Claim Verification 1,225 Claim, Evidence Label Macro-F1
Check-COVID Wang et al. (2023a) Scientific Claim Verification 1,504 Claim Evidence Accuracy, Precision, Recall, Macro-F1
HealthFC Vladika et al. (2024) Scientific Claim Verification 750 Claim, Evidence Label Precision, Recall, F1-Macro
FACTKG Kim et al. (2023) Scientific Claim Verification 108,000 Claim, Evidence Label Accuracy
BEAR-FACT Wührl et al. (2024a) Scientific Claim Verification 1,448 Claim, Evidence +Entity/Relation Information Label F1-Score
MINIF2F Zheng et al. (2022) Theorem Proving 488 Problem, Theorem Proof Pass Rate
FIMO Liu et al. (2023a) Theorem Proving 149 Problem, Theorem, statements Proof Pass Rate
LeanDojo Yang et al. (2023a) Theorem Proving 98,734 Problem, Theorem Proof R@k, MRR, Pass Rate
Lean-github Wu et al. (2024) Theorem Proving 28,597 Problem, Theorem Proof Accuracy, Pass Rate
TRIGO-real Xiong et al. (2023) Theorem Proving 427 Problem, Theorem Proof Pass Rate, Accuracy, EM@n
TRIGO-web Xiong et al. (2023) Theorem Proving 453 Problem, Theorem Proof Pass Rate, Accuracy, EM@n
TRIGO-gen Xiong et al. (2023) Theorem Proving - Problem, Theorem Proof Pass Rate, Accuracy, EM@n
CoqGym Yang and Deng (2019) Theorem Proving 71,000 Problem, Theorem Proof Success Rate
MLAgentBench Huang et al. (2024b) Experiment Validation 13 - - Competence, Efficiency
AAAR-1.0 Lou et al. (2024) Experiment Validation - Instance, Papers Design, Explanation S-F1, S-Precision, S-Recall + S-Match, ROUGE
TASKBENCH Shen et al. (2024) Experiment Validation 17,331 - - ROUGE, t-F1, v-F1 +Normalized Edit Distance
Spider2-V Cao et al. (2024a) Experiment Validation 494 Task Experiment Execution Success Rate
CORE-Bench Siegel et al. (2024) Experiment Validation 270 Task Requirements Experiment Result Accuracy
LAB-Bench Laurent et al. (2024) Experiment Validation 2400 Multiple-choice Question Answer Accuracy, Precision, Coverage
PaperBench Starace et al. (2025) Experiment Validation 20 Paper, Additional Information Code Replication Score
Hypothesis SUPER Bogin et al. (2024) Experiment Validation 801 Task Requirements - Accuracy, Landmark-Based Evaluation
Validation ScienceAgentBench Chen et al. (2025) Experiment Validation 102 Task Instruction, Dataset Information +Expert-Provided Knowledge Program Valid Execution Rate, Success Rate, CodeBERTScore, API Cost
SciCap+ Yang et al. (2023b) Manuscript Writing 414,000 Figure, OCR tokens + Mention Paragraph Caption BLEU, ROUGE, METEOR + CIDEr, SPICE
AAN Corpus Radev et al. (2013) Manuscript Writing - - - -
SciSummNet Yasunaga et al. (2019) Manuscript Writing 1,000 Paper,Citation Sentence Summary ROUGE
CiteBench Funkquist et al. (2023) Manuscript Writing 358,765 Cited Papers, Context Citation Text ROUGE, BERTScore
ALCE Gao et al. (2023b) Manuscript Writing 3,000 Question Answer with Citations Recall, Precision
GCite Wang et al. (2022b) Manuscript Writing 2,500 Citing/Cited Paper Citation Text BLEU, ROUGE
ARXIVEDITS Jiang et al. (2022b) Manuscript Writing 1,000 Sentence Pairs Sentence, Intent Precision,Recall,F1-score
CASIMIR Jourdan et al. (2024) Manuscript Writing 15,646 Original Sentence Revised Sentence Exact-match (EM),SARI, BLEU, + ROUGE-L,Bertscore
ParaRev Jourdan et al. (2025) Manuscript Writing 48,203 Original Paragraph Revised Paragraph ROUGE-L,SARI + BertScore
SCHOLAWRITE Wang et al. (2025) Manuscript Writing 62,000 Before-text Writing Intention, After-text F1-score, Lexical Diversity, Topic Consistency, Intention Coverage
MReD Shen et al. (2022) Peer Review 7,089 Reviews Meta-Review ROUGE
ORSUMZeng et al. (2024) Peer Review 15,062 Reviews Meta-Review ROUGE-L, BERTScore, FACTCC + SummaC, DiscoScore
PeerRead v1 Kang et al. (2018) Peer Review 107,000 Reviews Accept/Reject Accuracy
NLPeer Dycke et al. (2023) Peer Review 5,000 Reviews,Paper Review Score, Connection, + Review Category MRSE, F1-macro + Precision, Recall
AMPERE Hua et al. (2019) Peer Review 400 Review Review with Type Precision, Recall, F1-score
MOPRD Lin et al. (2023b) Peer Review 6,578 Reviews,Paper Editorial Decision, Review, + Meta-Review, Author Rebuttal ROUGE, BARTScore
ARIES D’Arcy et al. (2024b) Peer Review 1,720 Review Comment, Edits Comment-Edit Pairs Precision, Recall, F1-score
ASAP-Review Yuan et al. (2022) Peer Review - Paper Review Aspect Coverage, Aspect Recall, +Semantic Equivalence +Human: Recommendation Accuracy(RAcc), +Informativeness(Info),Aspect-level, +Constructiveness(ACon) and Summary accuracy
ReviewMT Tan et al. (2024) Peer Review 26,841 Paper Review Dialogue ROUGE,BLEU,METEOR
Manuscript ReAct Choudhary et al. (2021) Peer Review 6,250 Review Classification of Review Accuracy
Publication PEERSUM Li et al. (2023) Peer Review - Reviews Meta-Review ROUGE,BERTScore,UniEval,ACC

### v4

@@ 7.1 Integration of Diverse Research Tasks
The research process is an integrated pipeline of interdependent stages. Paper recommendation and literature review provide an AI tool with a field overview and relevant works, ensuring that hypothesis generation is informed and of higher quality. Hypothesis validation assesses feasibility both logically and practically, with results feeding back to refine the hypothesis (penades2025ai). In manuscript writing, validated hypotheses and prior outputs serve as key inputs. Peer review evaluates the manuscript and offers feedback across modules, enabling the hypothesis generator to adjust content accordingly (2408ChrisLuThe). In addition, combinations can also be made between some small fields, for instance, meta-review generation could be integrated with scientific claim verification, experiment verification could be linked with hypothesis formulation (jansen2025codescientist; 2501YuanJiakangDolphin; 2411ZijunLiuAIGS), and research paper recommendation systems could be connected with manuscript writing processes (2306NianlongGuSciLit). Furthermore, some studies have begun to emphasize the development of systems capable of covering multiple stages of the research process (2406PeterAJansenDiscoveryWorld; 2411YixuanWengCycleResearcher; 2412HaofeiYuResearchTown).
@@ 7.2 Integration with Reasoning-Oriented Language Models
Research is a process that places a significant emphasis on logic and reasoning. Theorem proving serves as a subtask within logical reasoning (2404ZhaoyuLiA), while hypothesis generation is widely recognized as the primary form of reasoning employed by scientists when observing the world and proposing hypothesis to explain these observations (2309ZonglinYangLarge). Experiment verification, in turn, demands a high degree of planning capability from models (2402SubbaraoKambhampatiPosition). Recent advances in reasoning-oriented language models, such as OpenAI-o1 (2412AaronJaechOpenAIo1) and DeepSeek-R1 (2501deepseekAIDeepseek-r1), have substantially enhanced the reasoning abilities of these models. Consequently, we posit that integrating reasoning language models with reasoning tasks is a promising future direction. This prediction was validated by experiments conducted by  2501SchmidgallAgent using o1-Preview.
@@
8 Ethical Considerations
AI has demonstrated significant potential in enhancing productivity by mitigating human limitations, thereby motivating increased investigation into its capacity to accelerate the research process  (2024MesseriArtificial). Nevertheless, its integration into scientific workflows introduces a range of ethical concerns  (2306BenediktFecherFriend; 2304MeredithScientists), including algorithmic biases, data privacy issues, risks of plagiarism, and the broader implications of AI-generated content for research communities. In this work, we examine these ethical challenges across the key stages of the research lifecycle: hypothesis formulation, validation, and publication.
During the hypothesis formulation stage, research paper recommendation systems and literature reviews are commonly employed; however, they often suffer from limitations that can lead to the formation of information bubbles and restrict exposure to diverse viewpoints. Furthermore, these systems tend to reinforce recognition disparities between prominent and lesser-known researchers and may inadvertently contribute to the dissemination of misinformation  (Polonioli21; 2402FranciscoBolaArtificial). To address these biases, recommendation algorithms can be enhanced by emphasizing content-based rather than author-based recommendations and by incorporating robust evaluation mechanisms to strengthen the credibility of suggested materials.
In contrast, AI-driven hypothesis generation presents more pronounced ethical challenges. First, the attribution of intellectual property rights and authorship for AI-generated hypotheses remains ambiguous  (2402BodhisattwaPrasadMajumder). Additionally, the widespread generation of low-quality content poses a risk of diluting the integrity of the academic landscape (2410XiangHuNova), while the potential misuse of such technologies for illicit purposes cannot be overlooked  (2409ChengleiSiCan). Addressing these concerns necessitates the development of robust accountability frameworks, the assignment of clear responsibility for AI-generated outputs to researchers, and the establishment of appropriate legal and regulatory mechanisms.
During the hypothesis validation phase, automated systems for scientific fact-checking remain underdeveloped. This limitation may be exploited by malicious actors to create advanced misinformation generators capable of circumventing existing fact-checking tools  (2112DavidWaddenMultiVerS). Likewise, in the context of experimental validation, there is a risk of unethical or legally questionable experiments being designed  (2502TransformingSteffen). These concerns underscore the need for continued research into model safety.
During the manuscript publication stage, several challenges remain. Text generated by AI models may carry a risk of plagiarism  (23SalvagnoCan; 2502TarunGuptaAll), while AI-assisted peer reviews often offer vague feedback and exhibit inherent biases  (2309LaurieA; 24IddoDroriHuman; 2502PataranutapornCan). To address these issues, the development of robust detection methods is essential. However, current detection tools are still in the early stages of maturity  (2502TarunGuptaAll).
@@ 9 Conclusion
We have made our best effort, but there may still be some limitations. Due to space constraints, we provide only concise summaries of each method without detailed technical elaboration. Given the rapid progress in AI and the expanding research landscape, we primarily focus on works published after 2022, with earlier studies receiving less attention. To emphasize areas that closely mimic the human research process, some topics are excluded from the main text but briefly discussed in Appendix §A. Moreover, as AI for Research is still an emerging field, the lack of standardized benchmarks and evaluation metrics hinders direct comparison. Nonetheless, we offer a comparative analysis of representative methods across domains using attribute graphs in Appendix §C.
Acknowledgments
Xiachong Feng and Xiaocheng Feng are the co-corresponding authors of this work. We thank the anonymous reviewers for their insightful comments. This work was supported by the National Natural Science Foundation of China (NSFC) (grant 62276078, U22B2059), the Key R&D Program of Heilongjiang via grant 2022ZX01A32, and the Fundamental Research Funds for the Central Universities ( XNJKKGYDJ2024013 ).